OpenStudy (angelface236):

x + 4/7 =9/5 x=63, x=56/12, 12/56, x= 43/35, x= 35/63, x=35/43

OpenStudy (redheadangel):

i would start by changing the fractions to decimals, it's easier to put into the calculator

OpenStudy (redheadangel):

so... x+.57=1.8

OpenStudy (redheadangel):

then you just subtract .57 from 1.8

OpenStudy (redheadangel):

which gets you 3.9

OpenStudy (mathmale):

Be careful: you are rounding off your decimals, which introduces imprecision. For precise answers, use the LCD to combine fractions with different original denominators. The LCD of x + 4/7 =9/5 is (7)(5)=35. Convert 4/7 and 9/5 both to fractions with 35 as their denominators.